part 1Required information [The following information applies to the questions displayed below.) Aspen Ski Resorts has 150 employeepart 2Required information [The following information applies to the questions displayed below.) Aspen Ski Resorts has 150 employeepart 3Required information [The following information applies to the questions displayed below.) Aspen Ski Resorts has 150 employee
Answer #1
1
Total salary expense 114000 =150*40*19
Federal Income tax 17100 =114000*15%
State Income tax 5700 =114000*5%
FICA tax 8721 =114000*7.65%
Total withholdings 31521
Actual direct deposit 82479 =114000-31521
2
FICA tax 8721 =114000*7.65%
Unemployment tax 7068 =114000*6.2%
Total payroll tax expense 15789
3
NO accounting entry required